# Segre forms of singular metrics on vector bundles and Lelong numbers

### v1

@@ 6. Segre forms and endomorphisms of qas bundles
Let $h_{0}^{\prime}$ denote another smooth reference metric. By (6.6) we see that ${\bf s}(E,h,h_{0}^{\prime})-{\bf s}(E,h,h_{0})$ is the image under $p_{*}$ of $(s(L_{E},e^{\psi_{0}^{\prime}})-s(L_{E},e^{\psi_{0}}))\wedge M^{\psi}$, which belongs to ${\mathcal{QZ}}_{0}({\mathbb{P}}(E))$. Now the first statement follows. The second one follows from Proposition 5.3 (ii) and (6.2).
It remains to prove (vii). By (6.4), it suffices to prove the result when $(E_{1},E_{2},k,\ell)$ is replaced by $(L_{E_{1}},L_{E_{2}},k+{\rm rank\,}E_{1}-1,\ell+{\rm rank\,}E_{2}-1)$. We may thus simply assume that ${\rm rank\,}E_{1}={\rm rank\,}E_{2}=1$. In that case, ${\bf s}_{k}={\bf s}_{1}^{k}$, so
$\big{(}{\bf s}_{k}(E_{1})\wedge{\bf s}_{\ell}(E_{2})-{\bf s}_{\ell}(E_{2})\wedge{\bf s}_{k}(E_{1})\big{)}\wedge\mu=({\bf s}_{1}(E_{1})^{k}\wedge{\bf s}_{1}(E_{2})^{\ell}-{\bf s}_{1}(E_{2})^{\ell}\wedge{\bf s}_{1}(E_{1})^{k})\wedge\mu.$
@@
$\mathbf{S}^{+}\wedge\big{(}{\bf s}_{1}(E_{1})\wedge{\bf s}_{1}(E_{2})-{\bf s}_{1}(E_{2})\wedge{\bf s}_{1}(E_{1})\big{)}\wedge\mathbf{S}^{-}\wedge\mu,$
@@
where $\mathbf{S}^{+}$ and $\mathbf{S}^{-}$ are products of ${\bf s}_{1}(E_{1})$ and ${\bf s}_{1}(E_{2})$. Let $Z=Z_{E_{1}}\cap Z_{E_{2}}$. It suffices to prove that each such term is of the form $dd^{c}w$, where $w$ is a current that is the push-forward of a current on $Z$. Since $\mathbf{S}^{-}\wedge\mu$ is in ${\mathcal{QZ}}$, we obtain by Lemma 6.7 below a current $w_{0}$, which is the push-forward of a current on $Z$ such that
$\big{(}{\bf s}_{1}(E_{1})\wedge{\bf s}_{1}(E_{2})-{\bf s}_{1}(E_{2})\wedge{\bf s}_{1}(E_{1})\big{)}\wedge\mathbf{S}^{-}\wedge\mu=dd^{c}w_{0},$
@@
where $w_{0}$ is current that is the push-forward of a current on $Z$. Note that $w_{0}$ has support on $Z$, and both ${\bf s}_{1}(E_{1})$ and ${\bf s}_{1}(E_{2})$ act by multiplication with a closed smooth form on currents in ${\mathcal{QZ}}$ with support on $Z$, and hence $\mathbf{S}^{+}$ acts as multiplication with a closed smooth form $S^{+}_{0}$, so we may take $w=S^{+}_{0}\wedge w_{0}$. ∎
Lemma 6.7.
If $E_{1},E_{2}$ are qas line bundles, $Z=Z_{E_{1}}\cap Z_{E_{2}}$ denote the intersection of the unbounded loci of $E_{1}$ and $E_{2}$. For any $\mu\in{\mathcal{QZ}}(X)$, there exists a current $w$, which is the push-forward of a current on $Z$ such that
$({\bf s}_{1}(E_{1})\wedge{\bf s}_{1}(E_{2})-{\bf s}_{1}(E_{2})\wedge{\bf s}_{1}(E_{1}))\wedge\mu=dd^{c}w,$
Proof.
By assuming that $\mu=\tau_{*}\gamma$, $\tau:W\to X$, and where $\gamma$ is of the form (4.1), we may assume by (6.9), after replacing $X$ by $W$, that $\mu$ is of the form (4.1). Let $\psi_{j}$ denote the metric on $E_{j}$, for $j=1,2$, and let $\omega_{j}$ denote the first Segre form of the smooth reference metric on $E_{j}$. Either some of $\psi_{j}\equiv-\infty$, say it is the case for $\psi_{1}$. Then ${\bf s}_{1}(E_{1})$ acts by multiplication with the smooth form $\omega_{1}$, which commutes with ${\bf s}_{1}(E_{2})$, and we may take $w_{0}=0$. Hence, we may assume that $\psi_{j}\not\equiv-\infty$.
As in the beginning of Section 4.1, we may perform a principalization of the singularities of $\psi_{j}$, and after a decomposition, we may assume that
$dd^{c}\psi_{j}=[D_{j}]+[W_{j}]+\omega_{j}^{\prime},$
where $D_{j},W_{j}$ are divisors for $j=1,2$, $|D_{1}|=|D_{2}|$, and $W_{1},W_{2},D_{1}$ intersect pairwise properly.
A calculation yields that
$({\bf s}_{1}(E_{1})\wedge{\bf s}_{1}(E_{2})-{\bf s}_{1}(E_{2})\wedge{\bf s}_{1}(E_{1}))\wedge\mu\\ =(([D_{1}]+[W_{1}]+\omega_{1}^{\prime})\wedge([W_{2}]+\omega_{2}^{\prime})+\omega_{1}\wedge[D_{2}]-([D_{2}]+[W_{2}]+\omega_{2}^{\prime})\wedge([W_{1}]+\omega_{1}^{\prime})+\omega_{2}\wedge[D_{1}])\wedge\mu\\ =([D_{1}]\wedge([W_{2}]+\omega_{2}^{\prime}-\omega_{2})-[D_{2}]\wedge([W_{1}]+\omega_{1}^{\prime}-\omega_{1}))\wedge\mu.$
We may choose qas metrics $h_{j}$ on ${\mathcal{O}}(D_{j})$ such that $\omega_{j}-[W_{j}]-\omega_{j}^{\prime}=s_{1}({\mathcal{O}}(D_{j}),h_{j})$, where the unbounded locus of $h_{j}$ equals $|W_{j}|$. Decompose $[D_{j}]=\sum d_{j,k}[D^{k}]$ for irreducible divisors $D_{k}$, and choose smooth reference metrics on each ${\mathcal{O}}(D^{k})$ with corresponding first Segre forms $\omega_{D^{k}}$. Then there is a qas function $u_{j}$ such that
$dd^{c}u_{j}=[W_{j}]+\omega_{j}^{\prime}+\sum_{k}d_{j,k}\omega_{D_{k}}-\omega_{j},$
and the set where $u_{j}$ is not locally bounded equals $|W_{j}|$. We may then take
$w=(u_{1}[D_{2}]-u_{2}[D_{1}])\wedge\mu,$
cf. the proof of Lemma 4.7. Note that indeed $w$ is the push-forward of a current of order 00 on $|D_{1}|=|D_{2}|=Z$. ∎
Proof of Theorem 1.4.
Items (i) to (v) are covered by Theorem 6.5 with $\mu=1$.
We now prove that these numbers are positive. First assume that $\psi$ is psh. In a small open neighborhood ${\mathcal{U}}$ of a given point on $X$, $E={\mathcal{U}}\times{\mathbb{C}}^{r}_{\alpha}$, and therefore $p^{-1}{\mathcal{U}}={\mathcal{U}}\times{\mathbb{P}}({\mathbb{C}}^{r}_{\alpha})$. Assume we choose a trivial metric $h_{0}$ on $E$, which implies that $\omega_{\alpha}=s_{1}(L_{E},h_{0})$ is positive. Assume now that $\psi$ is psh. In particular this holds if the metric is inherited from a morphism $g\colon E\to F$ and $F$ has a trivial metric, since then $\psi=\log|g(x)\alpha|^{2}$, which is psh. In that case, $s(L_{E},e^{\psi},e^{\psi_{0}})$ will be a positive current, and hence, so will $s(E,h,h_{0})=p_{*}(s(L_{E},e^{\psi},e^{\psi_{0}}))$. Furthermore, all components in the decomposition $s(E,h,h_{0})=s^{\prime}(E,h)+S^{E,h}+N^{E,h,h_{0}}$ will be positive as well, and in particular, all their components will have non-negative Lelong numbers.
By Lemma 3.7 any qas metric $h$ on $E$ is locally equivalent to a psh metric as above, and therefore all the concerned currents for $E$ with the metric $h$ have non-negative multiplicities. ∎
Remark 6.8.
The reference to Lemma 3.7 in last argument means that for the positivity of the multiplicities in part (vi), we need $E$ to be qas and not only that $L_{E}$ is qas, cf. Remark 3.9. The remainder of the theorem holds assuming only that $L_{E}$ is qas. ∎
Remark 6.9.
Almost all of Theorem 1.4 and Theorem 6.5 has natural extensions to ${\mathbb{Q}}$-qas or ${\mathbb{R}}$-qas metrics. Indeed, if $h$ is ${\mathbb{Q}}$-qas or ${\mathbb{R}}$-qas, it follows readily that ${\bf s}(E)$ yields an endomorphism of ${\mathcal{QZ}}_{\mathbb{Q}}(X)$ or ${\mathcal{QZ}}_{\mathbb{R}}(X)$, cf. Remark 4.11. In that case, the multiplicities referred to in the version of Theorem 1.4 would be non-negative rational or real numbers, cf. Remark 4.22. ∎
Example 6.10.
With the notation in the proof of Lemma 4.12, i.e., with $\mu=\tau_{*}\gamma$, we have
where the product on the right takes place on ${\mathbb{P}}(\tau^{*}E)$ and $\Pi=p\circ\tau^{\prime}=\tau\circ p^{\prime}$. The order of the factors in the product on ${\mathbb{P}}(\tau^{*}E)$ is important in general. To see (6.17) just notice that
$p_{*}([dd^{c}\psi]_{\omega}^{k+r-1})\wedge\tau^{\prime}_{*}(p^{\prime})^{*}\gamma)=p_{*}\tau^{\prime}_{*}([dd^{c}\psi]_{\omega}^{k+r-1})\wedge(p^{\prime})^{*}\gamma),$
and this holds because $\Gamma=(p^{\prime})^{*}\gamma$ is a product as in (4.1) and from (5.2) therefore
$[dd^{c}\psi]^{k+r-1}_{\omega}\wedge\tau^{\prime}_{*}\Gamma=\tau^{\prime}_{*}\big{(}[dd^{c}(\tau^{\prime})^{*}\psi]^{k+r-1}_{(\tau^{\prime})^{*}\omega}\wedge\Gamma\big{)}.$
Notice that
${\mathbb{P}}(\tau^{*}E)={\mathbb{P}}(E)\times_{X}W.$
Consider now (4.15) with $W={\mathbb{P}}(E)$ and $\tau=p$. Then we get the fiber diagram
(6.18) $\begin{array}[c]{cccc}Y&\stackrel{{\scriptstyle\pi_{1}}}{{\longrightarrow}}&{\mathbb{P}}(E_{1})\\ \downarrow\scriptstyle{\pi_{2}}&&\downarrow\scriptstyle{p_{1}}\\ {\mathbb{P}}(E_{2})&\stackrel{{\scriptstyle p_{2}}}{{\longrightarrow}}&X\end{array}$
with
(6.19) $Y={\mathbb{P}}(E_{1})\times_{X}{\mathbb{P}}(E_{2}),\quad\Pi=p_{1}\circ\pi_{1}=p_{1}\circ\pi_{2}\colon Y\to X,$
where $E_{1}$ and $E_{2}$ are distinct copies of $E\to X$. It follows from (6.2) that
${\bf s}_{\ell_{2}}(E,h,h_{0})\wedge s_{\ell_{1}}(E,h,h_{0})=\Pi_{*}\big{(}[dd^{c}\pi_{2}^{*}\psi_{2}]^{\ell_{2}+r-1}_{\pi_{2}^{*}\omega_{2}}\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r-1}_{\pi_{1}^{*}\omega_{1}}\big{)}=\\ (p_{2})_{*}([dd^{c}\psi_{2}]^{\ell_{2}+r-1}_{\omega_{2}}\wedge(\pi_{2})_{*}[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r-1}_{\pi_{1}^{*}\omega_{1}})=(p_{2})_{*}([dd^{c}\psi_{2}]^{\ell_{2}+r-1}_{\omega_{2}}\wedge p_{2}^{*}s_{\ell_{1}}(E,h,h_{0})),$
where $e^{\psi_{j}}$ and $e^{\psi_{j,0}}$ denote the metrics on $L_{E_{j}}$ induced by $h$ and $h_{0}$, and $\omega_{j}=s_{1}(L_{E_{j}},e^{\psi_{j,0}})$. Consider now $\mu=\tau_{*}\gamma$, where $\tau:W\to X$ and $\gamma\in{\mathcal{QZ}}(W)$. We may just as well consider $E_{1},\ldots,E_{k}$ to be $k$ copies of $E\to X$, $Y=W\times_{X}{\mathbb{P}}(E_{1})\times_{X}\cdots\times_{X}{\mathbb{P}}(E_{k})$, $\pi_{j}\colon Y\to{\mathbb{P}}(E_{j})$ and $\pi_{0}\colon Y\to W$ the natural projections. Then, similarly to above,
(6.20) ${\bf s}_{\ell_{k}}(E,h,h_{0})\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h,h_{0})\wedge\mu=\Pi_{*}([dd^{c}\pi_{k}^{*}\psi_{k}]_{\pi_{k}^{*}\omega_{k}}^{\ell_{k}+r-1}\wedge\cdots\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]_{\pi_{1}^{*}\omega_{1}}^{\ell_{1}+r-1}\wedge\pi_{0}^{*}\gamma\big{)}.$
Proposition 6.11.
Assume that $h$ is a qas metric on $E$ with empty unbounded locus. Let $h_{\epsilon}$ be the local regularization of $h$ from Proposition 6.4. If $\mu$ is in ${\mathcal{QZ}}$, then
$s_{\ell_{k}}(E,h_{\epsilon})\wedge\cdots\wedge s_{\ell_{1}}(E,h_{\epsilon})\wedge\mu\to{\bf s}_{\ell_{k}}(E,h)\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h)\wedge\mu.$
Proof.
Using notation from the previous example, and arguing as in the beginning of the proof of Proposition 6.4, it follows that if $\mu$ is in ${\mathcal{QZ}}$, then
${\bf s}_{\ell_{k}}(E,h)\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h)\wedge\mu=\Pi_{*}((dd^{c}\pi_{k}^{*}\psi_{k})^{\ell_{k}+r-1}\wedge\cdots\wedge(dd^{c}\pi_{1}^{*}\psi_{1})^{\ell_{1}+r-1}\wedge\pi_{0}^{*}\gamma\big{)}.$
Let $e^{\psi_{j,{\epsilon}}}$ denote the metrics corresponding to $h_{\epsilon}$. Then,
$s_{\ell_{k}}(E,h_{\epsilon})\wedge\cdots\wedge s_{\ell_{1}}(E,h_{\epsilon})\wedge\mu=\Pi_{*}((dd^{c}\pi_{k}^{*}\psi_{k,{\epsilon}})^{\ell_{k}+r-1}\wedge\cdots\wedge(dd^{c}\pi_{1}^{*}\psi_{1,{\epsilon}})^{\ell_{1}+r-1}\wedge\pi_{0}^{*}\gamma\big{)}.$
Letting ${\epsilon}\to 0$ yields the desired statement, just as in the proof of Proposition 6.4, see, e.g., [13, Theorem III.3.7]. ∎

### v2

@@ 6.2. The case with a general qas metric $h$
We now consider (vi). Let $h_{0}^{\prime}$ denote another smooth reference metric. By (6.6) we see that ${\bf s}(E,h,h_{0}^{\prime})-{\bf s}(E,h,h_{0})$ is the image under $p_{*}$ of $(s(L_{E},e^{\psi_{0}^{\prime}})-s(L_{E},e^{\psi_{0}}))\wedge M^{\psi}$, which belongs to ${\mathcal{QZ}}_{0}({\mathbb{P}}(E))$. Now the first statement follows. The second one follows from Proposition 5.2 (ii) and (6.2). Thus (vi) is proved. The proof of (vii) is postponed to Section 6.3. ∎
Proof of Theorem 1.3.
Items (i) to (v) are essentially covered by Theorem 6.7 with $\mu=1$. In the decomposition (1.4) in (iii), one must have that $M^{E,h,h_{0}}={\bf 1}_{Z}s(E,h,h_{0})$. By (4.6), $M^{h,h,h_{0}}$ belongs to ${\mathcal{QZ}}(X)$, and in particular, it is closed. By Theorem 6.5 (i), $s^{\prime}(E,h)={\bf 1}_{X\setminus Z}s(E,h,h_{0})$ is independent of $h_{0}$, and it coincides with the Segre form of $h$ on $X\setminus Z$ if $h$ is smooth there.
We now prove that these numbers are non-negative. In a small open neighborhood $\mathcal{U}$} of a given point on $X$, $E={\mathcal{U}}\times{\mathbb{C}}^{r}_{\alpha}$, and therefore $p^{-1}{\mathcal{U}}={\mathcal{U}}\times{\mathbb{P}}({\mathbb{C}}^{r}_{\alpha})$. If we choose a trivial metric $h_{0}$ on $E$, then $\omega_{\alpha}=s_{1}(L_{E},h_{0})$ is positive. Assume that the metric is inherited from a morphism $g\colon E\to F$ and $F$ has a trivial metric. Then $\psi=\log|g(x)\alpha|^{2}$ is psh. In that case, $s(L_{E},e^{\psi},e^{\psi_{0}})$ is a positive current, and hence, so is $s(E,h,h_{0})=p_{*}(s(L_{E},e^{\psi},e^{\psi_{0}}))$. Furthermore, each of the three terms in the decomposition $s(E,h,h_{0})=s^{\prime}(E,h)+S^{E,h}+N^{E,h,h_{0}}$ is positive as well, and in particular, all their components will have non-negative Lelong numbers. By Lemma 3.7 any qas metric $h$ on $E$ is locally equivalent to a psh metric as above. Thus (vi) is proved. ∎
Remark 6.9.
The proof of the positivity of the multiplicities refers to Lemma 3.7 which in turn requires that $E$ is qas and not only that $L_{E}$ is qas, cf. Remark 3.9. The remainder of the theorem holds assuming only that $L_{E}$ is qas. ∎
Remark 6.10.
Almost all of Theorem 1.3 and Theorem 6.7 have natural extensions to $\mathbb{Q}$}-qas or $\mathbb{R}$}-qas metrics. Indeed, if $h$ is $\mathbb{Q}$}-qas or $\mathbb{R}$}-qas, it follows readily that ${\bf s}(E,h,h_{0})$ yields an endomorphism on ${\mathcal{QZ}}_{\mathbb{Q}}(X)$ or ${\mathcal{QZ}}_{\mathbb{R}}(X)$, cf. Remark 4.11. In that case, the multiplicities referred to in the Theorem 1.3 would be non-negative rational or real numbers, cf. Remark 4.21. ∎
6.3. Composition of Segre endomorphisms
Example 6.11.
With the notation in the proof of Lemma 4.12, and $\mu=\tau_{*}\gamma$, we have
where the product on the right takes place on ${\mathbb{P}}(\tau^{*}E)$ and $\Pi=p\circ\tau^{\prime}=\tau\circ p^{\prime}$. To see (6.29) just notice that
$p_{*}([dd^{c}\psi]_{\omega}^{k+r-1}\wedge\tau^{\prime}_{*}(p^{\prime})^{*}\gamma)=p_{*}\tau^{\prime}_{*}([dd^{c}(\tau^{\prime})^{*}\psi]_{(\tau^{\prime})^{*}\omega}^{k+r-1}\wedge(p^{\prime})^{*}\gamma),$
and this holds because $\Gamma=(p^{\prime})^{*}\gamma$ is a product as in (4.1), and from (5.2) therefore
$dd^{c}\psi]^{k+r-1}_{\omega}\wedge\tau^{\prime}_{*}\Gamma=\tau^{\prime}_{*}\big([dd^{c}(\tau^{\prime})^{*}\psi]^{k+r-1}_{(\tau^{\prime})^{*}\omega}\wedge\Gamma\big).$
Notice that ${\mathbb{P}}(\tau^{*}E)={\mathbb{P}}(E)\times_{X}W$. ∎
Let $(E_{1},h_{1})$ and $(E_{2},h_{2})$ be two qas bundles over $X$. Consider (4.16) with $E=E_{2}$, $W={\mathbb{P}}(E_{1})$ and $\tau=p_{1}$. We then get the fiber diagram
(6.30) $\begin{array}[c]{cccc}Y&\stackrel{{\scriptstyle\pi_{2}}}{{\longrightarrow}}&{\mathbb{P}}(E_{2})\\ \downarrow\scriptstyle{\pi_{1}}&&\downarrow\scriptstyle{p_{2}}\\ {\mathbb{P}}(E_{1})&\stackrel{{\scriptstyle p_{1}}}{{\longrightarrow}}&X\end{array}$
with
(6.31) $Y={\mathbb{P}}(E_{2})\times_{X}{\mathbb{P}}(E_{1}),\quad\Pi=p_{1}\circ\pi_{1}=p_{2}\circ\pi_{2}\colon Y\to X.$
Let $h_{10}$ and $h_{20}$ be smooth reference metrics on $E_{1}$ and $E_{2}$ of rank $r_{1}$ and $r_{2}$, respectively. In view of Example 6.11 and (6.2) we have
${\bf s}_{\ell_{2}}(E_{2},h_{2},h_{20})\wedge s_{\ell_{1}}(E_{1},h_{1},h_{10})=(p_{2})_{*}([dd^{c}\psi_{2}]^{\ell_{2}+r_{2}-1}_{\omega_{2}}\wedge p_{1}^{*}s_{\ell_{1}}(E,h_{1},h_{10}))=\\ (p_{2})_{*}([dd^{c}\psi_{2}]^{\ell_{2}+r_{2}-1}_{\omega_{2}}\wedge(\pi_{2})_{*}[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r_{1}-1}_{\pi_{1}^{*}\omega_{1}})=\Pi_{*}\big([dd^{c}\pi_{2}^{*}\psi_{2}]^{\ell_{2}+r_{2}-1}_{\pi_{2}^{*}\omega_{2}}\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r_{1}-1}_{\pi_{1}^{*}\omega_{1}}\big).$
where $e^{\psi_{j}}$ and $e^{\psi_{j0}}$ denote the metrics on $L_{E_{j}}$ induced by $h_{j}$ and $h_{j0}$, and $\omega_{j}=s_{1}(L_{E_{j}},e^{\psi_{j0}})$. In the same way
(6.32) ${\bf s}_{\ell_{2}}(E_{2},h_{2},h_{20})\wedge{\bf s}_{\ell_{1}}(E_{1},h_{1},h_{10})\wedge\mu=\Pi_{*}\big([dd^{c}\pi_{2}^{*}\psi_{2}]^{\ell_{2}+r_{2}-1}_{\pi_{2}^{*}\omega_{2}}\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]^{\ell_{1}+r_{1}-1}_{\pi_{1}^{*}\omega_{1}}\wedge\Pi^{*}\mu\big),$
Example 6.12.
Let $(E_{1},h_{1}),\ldots,(E_{k},h_{k})$ be disjoint copies of $(E,h)$. Then, similarly,
(6.33) ${\bf s}_{\ell_{k}}(E,h,h_{0})\wedge\cdots\wedge{\bf s}_{\ell_{1}}(E,h,h_{0})\wedge\mu=\Pi_{*}([dd^{c}\pi_{k}^{*}\psi_{k}]_{\pi_{k}^{*}\omega_{k}}^{\ell_{k}+r-1}\wedge\cdots\wedge[dd^{c}\pi_{1}^{*}\psi_{1}]_{\pi_{1}^{*}\omega_{1}}^{\ell_{1}+r-1}\wedge\Pi^{*}\mu\big).$
Lemma 6.13.
Assume that $L_{1},L_{2}$ are qas line bundles over a manifold $Y$, and let $Z=Z_{L_{1}}\cap Z_{L_{2}}$ denote the intersection of their degeneracy loci. Let $Y^{\prime}\subset\subset Y$. For any $\mu\in{\mathcal{QZ}}(Y^{\prime})$ there exists a current $w$ on $Y^{\prime}$, that is the push-forward of a current on $Z\cap Y^{\prime}$, such that
$({\bf s}_{1}(L_{1})\wedge{\bf s}_{1}(L_{2})-{\bf s}_{1}(L_{2})\wedge{\bf s}_{1}(L_{1}))\wedge\mu=dd^{c}w.$
For simplicity we write ${\bf s}(L_{j})$ rather than ${\bf s}(L_{j},e^{\psi_{j}},e^{\psi_{j0}})$.
Proof.
By taking $\mu=\tau_{*}\gamma$, $\tau\colon W\to Y$, where $\gamma$ is of the form (4.1), we may assume by (6.21), after replacing $Y$ by $W$, that $\mu$ is of the form (4.1). Let $\psi_{j}$ denote the metric on $L_{j}$, for $j=1,2$, and let $\omega_{j}$ denote the first Segre form of the smooth reference metric on $L_{j}$. Either some of $\psi_{j}\equiv$-$\infty$, say it is the case for $\psi_{1}$. Then ${\bf s}_{1}(L_{1})$ acts by multiplication with the smooth form $\omega_{1}$, which commutes with ${\bf s}_{1}(L_{2})$, and we may take $w_{0}=0$. Hence we can assume that $\psi_{j}\not\equiv$-$\infty$ for both $j$.
In view of Proposition 5.2 (iii) we can replace $Y$ by any modification of it. As in the beginning of Section 4.1 we may therefore assume that
$dd^{c}\psi_{j}=[D_{j}]+[W_{j}]+\nu_{j},$
where $D_{j},W_{j}$ are divisors for $j=1,2$, $|D_{1}|=|D_{2}|$, and all irreducible components of $W_{1},W_{2},D_{1}$ intersect properly. Notice that ${\bf s}_{1}(L_{1})\wedge[D_{2}]\wedge\mu=\omega_{1}\wedge[D_{2}]\wedge\mu$ whereas
${\bf s}_{1}(L_{1})\wedge([W_{2}]+\nu_{2})\wedge\mu=([D_{1}]+[W_{1}]+\nu_{1})\wedge([W_{2}]+\nu_{2})\wedge\mu,$
where the right hand side can be computed formally since the intersections that occur are proper. We therefore get
(6.34) $({\bf s}_{1}(L_{1})\wedge{\bf s}_{1}(L_{2})-{\bf s}_{1}(L_{2})\wedge{\bf s}_{1}(L_{1}))\wedge\mu=([D_{1}]\wedge([W_{2}]+\nu_{2}-\omega_{2})-[D_{2}]\wedge([W_{1}]+\nu_{1}-\omega_{1}))\wedge\mu.$
If we choose a smooth metric on ${\mathcal{O}}(D_{j})$ we get a smooth representative $\alpha_{j}$ of its Bott-Chern class. Since $W_{j}]+\nu_{j}-\omega_{j}$ defines the same class there is a global function $u_{j}$ such that
$dd^{c}u_{j}=[W_{j}]+\nu_{j}-\omega_{j}-\alpha_{j}.$
Locally $u_{j}$ is a potential to $\nu_{j}$ plus a potential to $W_{j}]$ plus a smooth function, and hence it is qas. It now follows, cf. the proof of Lemma 4.7, that
$dd^{c}(u_{1}[D_{2}]\wedge\mu)=[dd^{c}u_{1}]\wedge[D_{2}]\wedge\mu=([W_{1}]+\nu_{1}-\omega_{1}-\alpha_{1})\wedge[D_{2}]\wedge\mu.$
Thus $w=(u_{1}[D_{2}]-u_{2}[D_{1}])\wedge\mu$ is a form with support on $Z=|D_{1}|=|D_{2}|$ such that, $dd^{c}w$ is equal to (6.34) minus
(6.35) $(\alpha_{1}[D_{2}]-\alpha_{2}[D_{1}])\wedge\mu.$
It remains to find a $dd^{c}$-potential to (6.35) in $Y^{\prime}$ that is the push-forward of a current on $Z\cap Y^{\prime}$. Since $Y^{\prime}\subset\subset Y$, $Z\cap Y^{\prime}$ has a finite number of irreducible components $D^{\ell}$. There are integers $a_{j}^{\ell}$ such that $D_{j}=\sum_{\ell}a_{j}^{\ell}D^{\ell},\ j=1,2.$ If $\omega^{\ell}$ are smooth representatives of the Bott-Chern class of $D^{\ell}]$, then we can choose $\alpha_{j}=\sum_{\ell}a_{j}^{\ell}\omega^{\ell}$ on $Y^{\prime}$. Thus
$(\alpha_{1}[D_{2}]-\alpha_{2}[D_{1}])\wedge\mu=\sum_{\ell\neq k}a_{1}^{\ell}a_{2}^{k}(\omega^{\ell}[D^{k}]-\omega^{k}[D^{\ell}])\wedge\mu.$
If $v_{\ell}$ are functions such that $dd^{c}v_{\ell}=[D^{\ell}]-\omega^{\ell}$, then as before, suppressing $\wedge\mu$ in the notation,
$dd^{c}(v_{\ell}[D^{k}]-v_{k}[D^{\ell}])=\omega^{\ell}[D^{k}]-\omega^{k}[D^{\ell}]-([D^{\ell}]\wedge[D^{k}]-[D^{k}]\wedge[D^{\ell}])=\omega^{\ell}[D^{k}]-\omega^{k}[D^{\ell}]$
since the $D^{\ell}]\wedge[D^{k}]$ is a proper intersection and thus commuting. Thus $w=\sum a_{1}^{\ell}a_{2}^{k}(v_{\ell}[D^{k}]-v_{k}[D^{\ell}])$ is a $dd^{c}$-potential to (6.35) as desired. ∎
Proof of Theorem 6.7 (vii).
Let $Y={\mathbb{P}}(E_{2})\times_{X}{\mathbb{P}}(E_{1})$ be as above, let $L_{j}\to Y$ be the pullback to $Y$ of the tautological line bundle $L_{E_{j}}\to{\mathbb{P}}(E_{j})$ and let $Y^{\prime}$ be the analogous set but over $X^{\prime}\subset\subset X$. Then $Y^{\prime}\subset\subset Y$. For simplicity we write $Y$ rather than $Y^{\prime}$ in what follows. By (6.32) it is enough to prove the similar statement for $L_{j}\to Y$, cf. (6.4), and $\ell,k$ replaced by $k+r_{2}-1,\ell+r_{1}-1$. Since $L_{j}$ are line bundles, ${\bf s}_{k}(L_{j})={\bf s}_{1}^{k}(L_{j})$, and so
$\big({\bf s}_{k}(L_{1})\wedge{\bf s}_{\ell}(L_{2})-{\bf s}_{\ell}(L_{2})\wedge{\bf s}_{k}(L_{1})\big)\wedge\mu=({\bf s}_{1}(L_{1})^{k}\wedge{\bf s}_{1}(L_{2})^{\ell}-{\bf s}_{1}(L_{2})^{\ell}\wedge{\bf s}_{1}(L_{1})^{k})\wedge\mu.$
@@ 6.3. Composition of Segre endomorphisms
$\mathbf{S}^{+}\wedge\big({\bf s}_{1}(L_{1})\wedge{\bf s}_{1}(L_{2})-{\bf s}_{1}(L_{2})\wedge{\bf s}_{1}(L_{1})\big)\wedge\mathbf{S}^{-}\wedge\mu,$
@@
where $\mathbf{S}^{+}$ and $\mathbf{S}^{-}$ are products of ${\bf s}_{1}(L_{1})$ and ${\bf s}_{1}(L_{2})$. Let $Z=Z_{L_{1}}\cap Z_{L_{2}}$. It suffices to prove that each such term is of the form $dd^{c}w$, where $w$ is a current that is the push-forward of a current on $Z$. Since $\mathbf{S}^{-}\wedge\mu$ is in $\mathcal{QZ}$}, we obtain by Lemma 6.13 a current $w_{0}$, which is the push-forward of a current on $Z$, such that
$\big({\bf s}_{1}(L_{1})\wedge{\bf s}_{1}(L_{2})-{\bf s}_{1}(L_{2})\wedge{\bf s}_{1}(L_{1})\big)\wedge\mathbf{S}^{-}\wedge\mu=dd^{c}w_{0},$
@@
where $w_{0}$ is current that is the push-forward of a current on $Z$. Note that $w_{0}$ has support on $Z$, and both ${\bf s}_{1}(L_{1})$ and ${\bf s}_{1}(L_{2})$ act by multiplication with a closed smooth form on currents in $\mathcal{QZ}$} with support on $Z$, and hence $\mathbf{S}^{+}$ acts as multiplication with a closed smooth form $S^{+}_{0}$, so we may take $w=S^{+}_{0}\wedge w_{0}$. ∎